OpenStudy (anonymous):

Please help!! ;( Which is an equation in point-slope form for the given point and slope? Point: (5, 9); Slope: 2 A. y – 5 = 2(x + 9) B. y + 9 = 2(x – 5) C. y – 9 = 2(x + 5) D. y – 9 = 2(x – 5)

jimthompson5910 (jim_thompson5910):

Point slope form is \[\Large y - y_1 = m(x - x_1)\]

jimthompson5910 (jim_thompson5910):

m is the slope (x1,y1) is the point that the line goes through

OpenStudy (anonymous):

i think it is b

OpenStudy (jdoe0001):

\(\bf Point:\ (5, 9); Slope:\ 2\\ \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 5}}\quad ,&{\color{blue}{ 9}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= 2 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values }\\ \qquad \uparrow\\ \textit{point-slope form} \)

OpenStudy (anonymous):

thanks for the quick replies :). i get it now